A planet of mass M moves around the Sun along an ellipse so that its minimum distance from the Sun is equal to r and the maximum distance to R. Making use of Kepler s laws, find its period of revolution around the Sun.
1
Expert's answer
2020-10-19T13:25:18-0400

Solution

Semi major axis "(\frac{r+R}{2}) =a"

According to kepler law time period is given by

"T=\frac{2\pi a^{3/2}}{\sqrt{ \gamma m_s}}"

"T=\pi\sqrt{\frac{(r+R) ^3}{2\gamma m_s}}"

Where ms is mass of sun.